OpenStudy (anonymous):

Which system of linear equations shares the same line? A. 6x − 3y = 15 8x − 4y = 20 B. 6x − 2y = 2 x + y = C. 3x − 3y = 15 9x − 6y = 6 D. −4x + 4y = −4 2x + 2y = 2

OpenStudy (anonymous):

I'll give medals and become a fan

OpenStudy (anonymous):

Please help.

OpenStudy (anonymous):

waht grade are u in

OpenStudy (anonymous):

8

OpenStudy (anonymous):

noo idea

OpenStudy (anonymous):

ok thanks anyways

OpenStudy (anonymous):

@ParthKohli do you know?

OpenStudy (akashdeepdeb):

A. 6x − 3y = 15 8x − 4y = 20 If you see in A 3(2x - y) = 3(5) => 2x-y = 5 AND 4(2x-y) = 4(5) => 2x-y=5 So for these types of questions, spot which one has a common factor and see if there simplest forms are equal or not! Understood? :)

OpenStudy (anonymous):

oh!!!

OpenStudy (anonymous):

yes it is A they both come out to y=2x-5

OpenStudy (anonymous):

ok thank you all!